OpenStudy (itiaax):

Given that the sum of the angles A, B and C of a triangle is π radians, show that: a) sin A = sin(B+C) b) sin A + sin B + sin C = sin(A+B) + sin(B+C) + sin(A+C) Please show all working clearly. Medal will be awarded.

OpenStudy (anonymous):

\[A+B+C=\pi ,B+C=\pi-A\] \[\sin \left( B+C \right)=\sin \left( \pi-A \right)=\sin A\] b. \[A+B+C=\pi,A+B=\pi-C,B+C=\pi-A,A+C=\pi-B\] Solve as a.

OpenStudy (itiaax):

I'm not quite understanding b

OpenStudy (anonymous):

from a.you find sin(B+C),similarly find sin(C+A),And Sin (A+B) and then add three.

OpenStudy (itiaax):

Can you tell me how you moved from sin(pi-a) to sina?

OpenStudy (itiaax):

I am still not getting b :(

OpenStudy (jdoe0001):

did you get the 1st part yet? a)

OpenStudy (itiaax):

No I havent @jdoe0001

OpenStudy (anonymous):

\[\sin \left( \pi-A \right)=\sin \pi \cos A-\cos \pi \sin A=0\times \cos A-\left( -1 \right)\sin A=\sin A\]

OpenStudy (itiaax):

Oh, now I get that part

OpenStudy (jdoe0001):

ahemmm ok

OpenStudy (jdoe0001):

the idea being, the reference angle 

OpenStudy (itiaax):

Can you show it in steps ?

OpenStudy (jdoe0001):

hmmm for ... part a) ?

OpenStudy (itiaax):

For part b